# Statistical Inference for Temporal Difference Learning with Linear Function Approximation

### v1

@@ 2 Background
For notional simplicity, we further introduce the notation $\bm{\Sigma}$ as feature covariance matrix
$\bm{\Sigma}:=\mathbb{E}_{s\sim\mu}[\bm{\phi}(s)\bm{\phi}^{\top}(s)],$ (7)
and use $\lambda_{\Sigma}$ and $\lambda_{0}$ to denote its largest and smallest eigenvalue, namely
$\lambda_{0}:=\lambda_{\min}(\bm{\Sigma}),\quad\text{and}\quad\lambda_{\Sigma}:=\lambda_{\max}(\bm{\Sigma}).$ (8)
The Temporal Difference (TD) learning algorithm.
To estimate the optimal vector $\bm{\theta}^{*}$ given by (4) we use the observe data, consisting of $T$ triplets of (state, action, reward),
$\{(s_{0},a_{0},r_{0}),(s_{1},a_{1},r_{1}),\ldots(s_{T},a_{T},r_{T})\},$
where, for each $t=0,1,\ldots,T$, $a_{t}\sim\pi(\cdot\mid s_{t})$, $r_{t}=r(s_{t},a_{t})$ and $s_{t+1}\sim P(\cdot\mid s_{t},a_{t})$.
The TD learning algorithm is a stochastic approximation algorithm that iteratively approximates the solution to the linear system $\bm{A}\bm{\theta}=\bm{b}$. Specifically, given a sequence of pre-selected step size $\{\eta_{t}\}_{t\geq 1},$ TD proceeds using the updating rule
$\bm{\theta}_{t}$ $=\bm{\theta}_{t-1}-\eta_{t}(\bm{A}_{t}\bm{\theta}_{t-1}-\bm{b}_{t}),$ (9a)
where
$\bm{A}_{t}$ $:=\bm{\phi}(s_{t-1})\left(\bm{\phi}(s_{t-1})-\gamma\bm{\phi}(s_{t})\right)^{\top}$ (9b)
and $\text{and }\qquad\bm{b}_{t}$ $:=\bm{\phi}(s_{t-1})r_{t-1}.$ (9c)
After a total number of $T$ iterations, the final iteration $\bm{\theta}_{T}$ is returned as the estimator for $\bm{\theta}^{\star}$.
Polyak-Ruppert averaging.
The Polyak-Ruppert averaging (Polyak and Juditsky,, 1992; Ruppert,, 1998) is a stochastic approximation methodology known to reduce the variability and accelerate the convergence of algorithms by averaging across the iterates. In addition, with a proper choice of stepsize, the averaged iterates converge to a Gaussian distribution asymptotically, which enables valid statistical inference that holds with large samples. When applied to the TD algorithm, Polyak-Ruppert averaging returns an estimator of $\bm{\theta}^{*}$ of the form
$\overline{\bm{\theta}}_{T}$ $=\frac{1}{T}\sum_{t=1}^{T}\bm{\theta}_{t},$ (10)
with associated estimation error
TD with Polyak-Ruppert averaging has been shown to deliver significant improvements, both of computational and statistical nature. While there is a recent strand of non-asymptotic results established for stochastic approximation (e.g. Mou et al., (2021); Moulines and Bach, (2011); Nemirovski et al., (2009); Lan, (2012)), the current results on Berry-Esseen bounds tailored to policy evaluation with linear function approximation still remain inadequate. In the work, we derive improved consistency and Gaussian approximation rates for $\overline{\bm{\Delta}}_{T}$. It is important to mention that the choice of learning rates $\{\eta_{t}\}_{t\geq 0}$ in the update (9) plays a critical role in determining the statistical properties of both the original and the averaged TD estimator, including their convergence rates and asymptotic distributions. This work focuses on the typical choice — the polynomial-decaying stepsizes — where $\eta_{t}=\eta_{0}t^{-\alpha}$ for $\alpha\in(1/2,1)$ and $\eta_{0}$ satisfying
$\eta_{0}<{1}/{(2\lambda_{\Sigma})}.$ (12)
3 Main results
While in most real-world RL scenarios the state-action-reward tuples are generated following a Markov decision process, there are applications where $i.i.d.$ samples become available, such as in game playing and autonomous driving. Furthermore, it is illuminating to first understand the statistical properties of this simpler setting before considering the more complicated setting of Markov samples. In this paper, like in most of the RL literature, we will make this simplifying assumption. Specifically, we will analyze the performance of the TD learning algorithm with Polyak-Ruppert averaging assuming a sample consisting of $T$ $i.i.d.$ state-action-reward tuples of the form
$\{(s_{t},a_{t},s_{t}^{\prime},r_{t})\}_{t=1}^{T},$ (13)
where $(s_{1},\ldots,s_{T})\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mu$ and, for eact $t$, $a_{t}\sim\pi(s_{t}),r_{t}=r(s_{t},a_{t})$ and $s_{t}^{\prime}\sim P(\cdot\mid s_{t},a_{t})$. Correspondingly, $\bm{A}_{t}$ and $\bm{b}_{t}$ are defined as
$\bm{A}_{t}$ $=\bm{\phi}(s_{t})\left(\bm{\phi}(s_{t})-\gamma\bm{\phi}(s_{t}^{\prime})\right)^{\top},$
$\text{and}\qquad\bm{b}_{t}$ $=r_{t}\bm{\phi}(s_{t}).$
In this setting, with polynomial-decaying stepsizes, it has been shown (Polyak and Juditsky, (1992); Borkar, (2008); Fort, (2015)) that the estimation error satisfies
[…]
where the covariance matrix $\bm{\Lambda}^{\star}$ is given as
$\bm{\Lambda}^{\star}=\bm{A}^{-1}\bm{\Gamma}\bm{A}^{-\top},\quad\text{with}$ (15)
$\bm{\Gamma}=\mathbb{E}[(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}].$ (16)
@@ 3 Main results
which may serve as a comparison term. Our main goal is to demonstrate how the TD algorithm can be used to construct confidence regions and simultaneous confidence intervals for $\bm{\theta}^{*}$, the linear parameters of the best value function linear approximation (see (5)). Towards this goal, we shall establish a high-probability convergence guarantee for $\overline{\bm{\Delta}}_{T}$, develop high-dimensional Berry-Esseen bounds over the class of convex sets, and finally build confidence regions/intervals for $\bm{\theta}^{\star}$ using the estimator $\overline{\bm{\theta}}_{T}$. These topics will be addressed in order.
3.1 High-probability convergence guarantees
We start by controlling the convergence rate of the TD iterates with Polyak-Ruppert averaging in high probability. Before stating our main results, let us first introduce additional quantities
where $\bm{\Gamma}$ is defined in (16) an, for each $t$, we set
$\bm{Q}_{t}=\eta_{t}\sum_{j=t}^{T}\prod_{k=t+1}^{j}(\bm{I}-\eta_{k}\bm{A}).$ (19)
The matrix $\overline{\bm{\Lambda}}_{T}$ plays the role of a non-asymptotic covariance matrix for $\overline{\bm{\theta}}_{T}$.
@@ 3.1 High-probability convergence guarantees
Consider TD with Polyak-Ruppert averaging (9) with independent samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$ for $\alpha\in(\frac{1}{2},1)$. For any $\delta\in(0,1)$, there exists $\eta_{0}=\eta_{0}(\delta)>0$, such that
$\|\overline{\bm{\Delta}}_{T}\|_{2}$ $\lesssim\sqrt{\frac{\mathsf{Tr}(\overline{\bm{\Lambda}}_{T})}{T}\log\frac{d}{\delta}}+\widetilde{C}T^{-\frac{\alpha+1}{2}}\log\frac{T}{\delta}$
with probability at least $1-\delta$, where $\widetilde{C}$ is a problem-related quantity independent of $T$ and $\delta$.
The proof can be found in Appendix B.2, where the value of $\widetilde{C}$ is given in equation (75). Throughout the paper, for the sake of readability, we have opted to express unwieldy problem-dependent quantities that do not directly depend on $T$ in implicit form; their values can be tracked in the proofs. Accordingly, when writing our finite-sample bounds, we will first present the leading terms in $T$.
In a nutshell, Theorem 3.1 states that $\overline{\bm{\theta}}_{T}$ converges to $\bm{\theta}^{*}$ in $\ell_{2}$ error at a rate $O(\sqrt{{\mathsf{Tr}(\overline{\bm{\Lambda}}_{T})}/{T}})$ up to log factor. It is worth comparing the high-probability sample complexity bound of Theorem 3.1 with those in the literature. To the best of our knowledge, the tightest results were attained by Dalal et al., (2018) for inverse-decaying stepsizes and Li et al., 2024b ; Samsonov et al., (2023) for constant stepsize as summarized in Table 1. With polynomial-decaying stepsizes, Theorem 3.1 is the first of its kind to achieve tight high-probability convergence guarantees up to log factors. Another subtle but consequential difference concerns the restriction on the step size. For our bound to hold, the choice of the initial stepsize $\eta_{0}$ depends on $\delta$, and is required to be sufficiently small to satisfy $\eta_{0}\sqrt{\log{1}/(\delta\eta_{0}})\leq\widetilde{C}$, where $\widetilde{C}$ is independent of $T$ and $\delta$ but may depend on the other parameters of interest. Though not ideal, this condition is a considerable weakening of the stepsize requirements imposed by Samsonov et al., (2023) and Li et al., 2024b , which depend on both $\delta$ and $T$; see Table 1.
The bound of Theorem 3.1 can be made dependent on the asymptotic variance $\bm{\Lambda}^{\star}$ without affecting the rate. Indeed, as we will see below, the difference between $\overline{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$ converges at the rate of $O(T^{\alpha-1})$. Thus, the dominant terms in Theorem 3.1 match the convergence rate (17) implied by the asymptotic distribution up to a logarithm term.
@@ 3.2 High-dimensional Berry-Esseen bounds
Next we derive high-dimensional Berry-Esseen bounds for TD with Polyak-Ruppert averaging. We first characterize the convex distance between $\sqrt{T}\overline{\bm{\Delta}}_{T}$ and $\overline{\bm{\Lambda}}_{T}^{1/2}\bm{z}$, where $\bm{z}$ is a $d$-dimensional standard gaussian. The proof of this result is included in Appendix B.3.
@@
Consider the TD with Polyak-Ruppert averaging with independent samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$ for $\alpha\in(\frac{1}{2},1)$. Then,
$d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\overline{\bm{\Lambda}}_{T}^{1/2}\bm{z})\lesssim\sqrt{\frac{\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}}\mathsf{Tr}(\bm{\Gamma})\|\bm{\Gamma}^{-1}\|T^{-\frac{\alpha}{2}}+\widetilde{C}\;T^{-\frac{1}{2}}\log T,$
where $\widetilde{C}$ is a problem related quantity indepedent of $T$.
To relate this result with the classical CLT (as in (14)), we will use the following comparison bound for the total variation distance between $\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T})$ and $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$; see Appendix B.4 for the proof.
Theorem 3.3.
Let $\bm{\Lambda}^{\star}$ and $\overline{\bm{\Lambda}}_{T}$ be defined as in (15) and (18), respectively, with the stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$, for all $t$ and any $\alpha\in(\frac{1}{2},1)$. Then,
@@
$d_{\mathsf{C}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\leq d_{\mathsf{TV}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\lesssim\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma})}}{(1-\gamma)\lambda_{0}\eta_{0}}T^{\alpha-1}+\widetilde{C}\;T^{2\alpha-2},$
where $\widetilde{C}$ is a problem related quantity indepedent of $T$.
Comparison with existing results.
The latest and sharpest Berry-Esseen bounds for stochastic approximation under condition (12) are du to Samsonov et al., (2024). Specifically, the authors show that
@@
$d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},{\bm{\Lambda}}^{\star 1/2}\bm{z})\leq\widetilde{C}\;T^{\frac{\alpha-1}{2}},$
where $\widetilde{C}$ depend on other problem parameters. Upon choosing $\alpha=\frac{1}{2}$, they obtain a convergence rate of order $O(T^{-\frac{1}{4}})$. In contrast, the Berry-Esseen bounds of Theorem 3.2 and the comparison bound of Theorem 3.3 indicate that, selecting the optimal $\alpha=\frac{2}{3}$ leads to the improved rate
$d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\bm{\Lambda}^{\star 1/2}\bm{z})\leq\widetilde{C}\;T^{-\frac{1}{3}}.$
The proof of Theorem 3.2 relies on Theorem 3.1 of Shao and Zhang, (2021), a high-dimensional Berry-Esseen bound for non-linear statistics, which is also used by Samsonov et al., (2024) to establish their main result about Gaussian approximations for the TD error, Theorem 3.4 therein. We are able to improve their the rate of convergence from $O(T^{\frac{1}{2}-\alpha})$ to $O(T^{-\frac{\alpha}{2}})$ by carefully handling some of the quantities appearing in the bound of Theorem 3.1 of Shao and Zhang, (2021) [^2] Specifically, the term $I_{3}$ of Equation (79) in Appendix B.3.. This sharpening is non-trivial and may be of independent interest, as it may be potentially generalized to analyze stochastic gradient algorithms applied to non-linear functions (as of Theorem 3.4 of Shao and Zhang, (2021)). We leave this extension to future work.
@@ 3.3 Online estimation of the covariance matrix
Thus far, we have provided high-dimensional Berry-Esseen bounds to quantify the non-asymptotic discrepancy between $\overline{\bm{\Delta}}_{T}$ and its Gaussian limit. In order to invoke these results for constructing confidence region or confidence intervals for $\bm{\theta}^{\star}$, it is necessary to estimate the covariance matrix $\bm{\Lambda}^{\star}$. Two types of approaches have been proposed recently, namely plug-in estimation (Chen et al., (2020); Zhu et al., (2023); Li et al., (2022)) and bootstraping (White and White, (2010); Hanna et al., (2017); Hao et al., (2021); Ramprasad et al., (2022); Fang et al., (2018); Samsonov et al., (2024)); nontheless, the statistical and computational efficiency of these estimations can both be further enhanced.
In this section, we introduce a computationally feasible online plug-in estimator for $\bm{\Lambda}^{\star}$, and establish its non-asymptotic rate of convergence. We will use
$\overline{\bm{A}}_{T}=\frac{1}{T}\sum_{t=1}^{T}\bm{A}_{t};$ (20)
$\widehat{\bm{\Gamma}}_{T}=\frac{1}{T}\sum_{t=1}^{T}(\bm{A}_{t}\overline{\bm{\theta}}_{T}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\theta}}_{T}-\bm{b}_{t})^{\top},$ (21)
@@
Crucially, because of the linear nature of the problem, the estimator $\widehat{\bm{\Lambda}}_{T}$ can be evaluated in an online manner. Specifically, the basic properties of Kronecker products indicate
Since Algorithm 1 iterates from $t=0$ to $t=T-1$ in a single loop and does not store the samples for each iteration, it has a space complexity is $O(Td^{4})$ and a sapce complexity of $O(d^{4})$. Here, the $d^{4}$ term stems from calculating the Kronecker product $\bm{A}\otimes\bm{A}$. Not only is the estimator $\widehat{\bm{\Lambda}}_{T}$ computationally feasible, but it also enjoys good statistical properties. This is made precise by the following Gaussian comparison result, whose proof is in Appendix B.5, showing that the unknown asymptotic variance $\bm{\Lambda}^{\star}$ in the Berry-Esseen bound of Theorem 3.3 can be replaced by $\widehat{\bm{\Lambda}}_{T}$ without compromising the accuracy of the approximation.
@@
Consider the TD algorithm (9) with independent samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$ for all $t$ and any $\alpha\in(1/2,1)$, and let $\widehat{\bm{\Lambda}}_{T}$, $\bm{\Lambda}^{\star}$ be defined as in (22), (15) respectively. Then,
@@
Comparison with Chen et al., (2020).
In different settings, Chen et al., (2020) considered the problem of carrying out statistical inference for the true model parameters using stochastic gradient descent and assuming a strongly convex and smooth population loss function. In their work, a plug-in estimator of the form $\widehat{\bm{\Lambda}}_{T}^{\prime}=\overline{\bm{A}}^{-1}\widehat{\bm{\Gamma}}_{T}^{\prime}\overline{\bm{A}}^{\top}$, where
$\widehat{\bm{\Gamma}}_{T}^{\prime}:=\frac{1}{T}\sum_{t=1}^{T}(\bm{A}_{t}{\bm{\theta}}_{t}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}}_{t}-\bm{b}_{t})^{\top},$ (23)
is proposed, which differs slightly from ours. The authors proved that the estimation error of $\widehat{\bm{\Lambda}}_{T}^{\prime}$ is bounded in expectatin by
@@
where $\widetilde{C}$ depends on $\lambda_{0},\gamma,\eta_{0},\alpha,\bm{\theta}^{\star}$ and $d$. Comparing (24) with (25), we have tightened the bound from $O\left(T^{-\frac{\alpha}{2}}\right)$ to $O\left(T^{-\frac{1}{2}}\right)$, independent of $\alpha\in(1/2,1)$. This is a significant improvement, at least in our setting, not only because of the different metric but because a root-$T$ rate based on (24) would require $\alpha=1$, which is not permissible in our problem. It is noteworthy that the online algorithm 1 takes advantage of the linearity of the problem; in general, potentially non-linear settings, replacing $\bm{\theta}_{t}$ in (23) by $\overline{\bm{\theta}}_{T}$ in (21) would require storing all the samples throughout the iterative trajectory, resulting in a larger space complexity.
Comparison with Samsonov et al., (2024).
@@
Using a multiplier-bootstrap technique, Samsonov et al., (2024) constructs a random estimator based on the samples $\mathcal{S}_{T}:=\{s_{t},s_{t}^{\prime}\}_{t=1}^{T}$, denoted as $\overline{\bm{\theta}}_{T}^{(b)}$.Theorem 4 and 5 in Samsonov et al., (2024)) imply that
@@
$d_{\mathsf{C}}(\sqrt{T}(\overline{\bm{\theta}}_{T}^{(b)}-\bm{\theta}^{\star})|\mathcal{S}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\lesssim\frac{\|\bm{\Gamma}^{-1}\|}{[\lambda_{0}(1-\gamma)]^{\frac{3}{2}}}(2\|\bm{\theta}^{\star}\|_{2}+1)^{2}\frac{\log T}{T^{\frac{1}{4}}}$ (26)

### v3

@@ 3 Results
where the asymtoptic covariance matrix $\bm{\Lambda}^{\star}$ is given by
$\bm{\Lambda}^{\star}=\bm{A}^{-1}\bm{\Gamma}\bm{A}^{-\top},\quad\text{with}\quad\bm{\Gamma}=\mathbb{E}[(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}].$ (12)
@@
which may serve as a benchmark. Our main goal is to demonstrate how the TD algorithm can be used to construct non-asymptotic confidence regions and simultaneous confidence intervals for $\bm{\theta}^{*}$ in high dimensions. Towards this goal, in the subsequent sections we will establish a high-probability convergence guarantee for $\overline{\bm{\Delta}}_{T}$, develop high-dimensional Berry-Esseen bounds over the class of convex sets, and finally build confidence regions/intervals for $\bm{\theta}^{\star}$ using the estimator $\overline{\bm{\theta}}_{T}$.
3.1 High-probability bounds
We start with a high-dimensional, high-probability bound on the size of the TD estimation error.
@@ 3.1 High-probability bounds
Consider the TD algorithm with Polyak-Ruppert averaging (7), independent samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$ for $\alpha\in(\frac{1}{2},1)$. For any $\delta\in(0,1)$, there exists $\eta_{0}=\eta_{0}(\delta)>0$, such that
$\|\overline{\bm{\Delta}}_{T}\|_{2}$ $\lesssim\sqrt{\frac{\mathsf{Tr}(\bm{\Lambda}^{\star})}{T}\log\frac{1}{\delta}}+\tilde{C}T^{-\frac{\alpha+1}{2}}\log\frac{T}{\delta}$
with probability at least $1-\delta$, where $\tilde{C}$ is a problem-related quantity independent of $T$ and $\delta$.
The proof can be found in Appendix B.1, where the value of $\tilde{C}$ is given in equation (78). Throughout the paper, for the sake of readability, we have opted to express unwieldy problem-dependent quantities that do not directly depend on $T$ in implicit form; their values can be tracked in the proofs. Accordingly, when writing our finite-sample bounds, we will first present the leading terms in $T$.
In a nutshell, Theorem 3.1 states that $\overline{\bm{\theta}}_{T}$ converges to $\bm{\theta}^{*}$ (in Euclidean distance) at the rate of $O(\sqrt{{\mathsf{Tr}(\bm{\Lambda}^{\star})}/{T}})$ up to a log factor. It is worth comparing the high-probability sample complexity bound of Theorem 3.1 with those in the literature. To the best of our knowledge, the tightest results were attained by Dalal et al. (2018) for inverse-decaying stepsizes and Li et al. (2024); Samsonov et al. (2024) for constant stepsize as summarized in Table 1. Theorem 3.1 is the first result to achieve tight high-probability convergence guarantees up to log factors using polynomial-decaying stepsizes. Another subtle but consequential difference concerns the restriction on the step size. For our bound to hold, the choice of the initial stepsize $\eta_{0}$ depends on $\delta$; specifically, $\eta_{0}$ is required to be sufficiently small such that $\eta_{0}\sqrt{\log{1}/(\delta\eta_{0}})\leq\tilde{C}$, where $\tilde{C}$ is independent of $T$ and $\delta$ but depends on the other parameters of interest; its specific form is shown in (70) and (71) in the Appendix. Though not ideal, this condition is necessary for a Chernoff-type convergence guarantee; we refer to Appendix B.1.3 for a detailed discussion. Nonetheless, our stepsize requirements are considerably weaker than those imposed by Samsonov et al. (2024) and Li et al. (2024), which depend on both $\delta$ and $T$; see Table 1.
@@ 3.2 High-dimensional Berry-Esseen bounds
In our next result, we derive novel high-dimensional Berry-Esseen bounds for the TD learning estimator, certifying that, uniformly over the class of $d$-dimensional convex sets, $\sqrt{T}\overline{\bm{\Delta}}_{T}$ is close in distribution to a centered Gaussian random vector with the optimal asymptotic covariance (12), thus closely mimicking the classical, fixed-dimensional CLT guarantee in (11). We reiterate that, to the best of our knowledge, the resulting rates are the sharpest to date.
@@
Consider the TD algorithm with Polyak-Ruppert averaging (7), independent samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$ for $\alpha\in(\frac{1}{2},1)$ and $\eta_{0}<1/2$. Then,
@@
$d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))$ $\lesssim\sqrt{\frac{\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}}\mathsf{Tr}(\bm{\Gamma})\|\bm{\Gamma}^{-1}\|T^{-\frac{\alpha}{2}}+\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma})}}{(1-\gamma)\lambda_{0}\eta_{0}}T^{\alpha-1}$
$+\tilde{C}\;T^{-\frac{1}{2}}+\tilde{C}^{\prime}\;T^{2\alpha-2}$
where $\tilde{C},\tilde{C}^{\prime}$ are problem-related quantities independent of $T$.
On the final rate. By choosing $\alpha=2/3$, we obtain a rate in $T$ of order $O(T^{-1/3})$. In fact, in Theorem B.3 in the Appendix we establish the more refined result that the Berry-Esseen bound (in the convex distance) between the distribution of $\sqrt{T}\overline{\bm{\Delta}}_{T}$ and the Gaussian distribution with (approximately) matching first two moments is of order $\sqrt{\frac{\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}}\mathsf{Tr}(\bm{\Gamma})\|\bm{\Gamma}^{-1}\|T^{-\frac{\alpha}{2}}$. As $\alpha\to 1$, this rate approaches the optimal rate of $1/\sqrt{T}$. The additional term $\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma})}}{(1-\gamma)\lambda_{0}\eta_{0}}T^{\alpha-1}$ in the above bound, which is ultimately the reason for the final rate $O(T^{-1/3})$, originates from replacing an approximation to the finite sample variance of $\sqrt{T}\overline{\bm{\Delta}}_{T}$ with its asymptotic variance $\bm{\Lambda}^{\star}$.
Comparison with Samsonov et al. (2024). Prior to this work, the best Berry-Esseen bounds for stochastic approximation with polynomial-decaying stepsizes and Polyak-Ruppert averaging are those obtained by Samsonov et al. (2024), thus covering the settings considered here. Specifically, the authors show in their Theorem 3 that when $\eta_{0}\leq\min\{\lambda_{0}(1-\gamma),\frac{1-\gamma}{(1+\gamma)^{2}}\}$ and $\alpha\in[\frac{1}{2},1)$,
@@
$d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\lesssim O(T^{-\frac{\alpha}{2}}+T^{\frac{\alpha-1}{2}}).$
Choosing $\alpha=\frac{1}{2}$ leads to a convergence rate of order $O(T^{-\frac{1}{4}})$. In contrast, Theorem 3.2 shows that, by selecting $\alpha=\frac{2}{3}$, we obtain the faster rate $O(T^{-\frac{1}{3}})$. As shown in Appendix B.2.3, this rate is tight for $\alpha\in(\frac{2}{3},1)$. See also Figure 2 below, showing the output of a numerical experiment demonstrating that the choice of $\alpha=\frac{2}{3}$ leads to a smaller approximation error. Though worked out independently, the proofs of our Theorem 3.2 and of Theorem 3 of Samsonov et al. (2024) bear similarities, as they both rely on the general Berry-Esseen bounds for non-linear statistics developed by Shao and Zhang (2022). We obtain sharper rates by a more targeted analysis of the TD settings, while, in comparison, the results of Samsonov et al. (2024) apply more broadly to SA problems. Specifically we (i) adopt a different decomposition of the TD estimation error, (ii) handle in a more refined way certain terms appearing in the bound of Theorem 3.1 of Shao and Zhang (2022) [^2] Specifically, the term $I_{3}$ in Appendix B.2.1. and (iii) carry out a more specialized analysis of the properties of the matrices $\bm{A}_{t}$ and $\bm{A}$ (see, e.g., Lemma A.4 in the Appendix) originating from TD learning, which has allowed us to tighten problem-related coefficient, as well as to work with a substantially larger intial stepsize $\eta_{0}$. This non-trivial sharpening may be potentially generalized to analyze stochastic gradient algorithms applied to non-linear functions, as, e.g., in Theorem 3.4 of Shao and Zhang (2022).
@@ 3.3 Online estimation of the covariance matrix
In the last section, we obtained finite sample bounds on the convex distance between $\sqrt{T}\overline{\bm{\Delta}}_{T}$ and its Gaussian approximation, which features an unknown variance. To construct confidence regions for $\bm{\theta}^{\star}$, it is necessary to estimate the covariance matrix $\bm{\Lambda}^{\star}$. Two types of approaches have been recently proposed: plug-in estimation (Chen et al., 2020) and bootstraping (White and White, 2010; Hanna et al., 2017; Hao et al., 2021; Ramprasad et al., 2022; Zhu et al., 2023; Samsonov et al., 2024).
Here we propose to use a computationally feasible online plug-in estimator for $\bm{\Lambda}^{\star}$ with superior computational efficiency than existing methodologies and establish its non-asymptotic rate of convergence. We will use
$\overline{\bm{A}}_{T}=\frac{1}{T}\sum_{t=1}^{T}\bm{A}_{t},\quad\text{and}\quad\widehat{\bm{\Gamma}}_{T}=\frac{1}{T}\sum_{t=1}^{T}(\bm{A}_{t}\overline{\bm{\theta}}_{T}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\theta}}_{T}-\bm{b}_{t})^{\top},$ (14)
@@
Crucially, because of the linear nature of the problem, the estimator $\widehat{\bm{\Lambda}}_{T}$ can be evaluated in an online manner. See Algorithm 1 in Appendix D.1 for a detailed description of the procedure, which has time and space complexities of order $O(Td^{4})$ and $O(d^{4})$, respectively. The estimator $\widehat{\bm{\Lambda}}_{T}$ also enjoys good statistical properties. This is made precise by the following Gaussian comparison result, whose proof can be found in Appendix B.3, showing that the unknown asymptotic variance $\bm{\Lambda}^{\star}$ in the Berry-Esseen bound of Theorem B.4 can be replaced by $\widehat{\bm{\Lambda}}_{T}$ without compromising the accuracy of the approximation (and without affecting the rate in $T$).
@@
Consider the TD algorithm (7) with independent samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$ for all $t>0$, any $\alpha\in(1/2,1)$ and $\eta_{0}<1/2$ , and let $\widehat{\bm{\Lambda}}_{T}$, $\bm{\Lambda}^{\star}$ be defined as in (15), (12) respectively. Then,
@@
Comparison with Chen et al. (2020). In different settings, Chen et al. (2020) considered the problem of carrying out statistical inference for the true model parameters using stochastic gradient descent and assuming a strongly convex and smooth population loss function. In their work, they consider the slightly different plug-in estimator $\widehat{\bm{\Lambda}}_{T}^{\prime}=\overline{\bm{A}}_{T}^{-1}\widehat{\bm{\Gamma}}_{T}^{\prime}\overline{\bm{A}}_{T}^{\top}$, where $\widehat{\bm{\Gamma}}_{T}^{\prime}:=\frac{1}{T}\sum_{t=1}^{T}(\bm{A}_{t}{\bm{\theta}}_{t}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}}_{t}-\bm{b}_{t})^{\top}$. The authors proved that the estimation error of $\widehat{\bm{\Lambda}}_{T}^{\prime}$ is bounded in expectation as
@@
where $\widetilde{C}$ depends on $\lambda_{0},\gamma,\eta_{0},\alpha$ and $\bm{\theta}^{\star}$. Comparing (16) with (17), we have sharpened the rate from $O\left(T^{-\frac{\alpha}{2}}\right)$ to $O\left(T^{-\frac{1}{2}}\right)$, independent of the choice of $\alpha$.
@@
It is noteworthy that the online Algorithm 1 takes advantage of the linearity of the problem; in general, potentially non-linear settings, replacing $\bm{\theta}_{t}$ in $\widehat{\bm{\Gamma}}_{T}^{\prime}$ by $\overline{\bm{\theta}}_{T}$ in $\widehat{\bm{\Gamma}}_{T}$ would require storing all the samples throughout the iterative trajectory, resulting in a larger space complexity.
Comparison with Samsonov et al. (2024).
@@
Leveraging the multiplier-bootstrap technique proposed by Zhu et al. (2023), Samsonov et al. (2024) constructs a randomized estimator $\overline{\bm{\theta}}_{T}^{(b)}$ based on the samples $\mathcal{S}_{T}:=\{s_{t},s_{t}^{\prime}\}_{t=1}^{T}$. Their Theorems 4 and 5 imply that, in our notation,
@@
$d_{\mathsf{C}}(\sqrt{T}(\overline{\bm{\theta}}_{T}^{(b)}-\bm{\theta}^{\star})|\mathcal{S}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\lesssim\frac{\|\bm{\Gamma}^{-1}\|}{[\lambda_{0}(1-\gamma)]^{\frac{3}{2}}}(2\|\bm{\theta}^{\star}\|_{2}+1)^{2}\frac{\log T}{T^{\frac{1}{4}}}+o(T^{-\frac{1}{4}})$ (18)